Which label describes the following? 3x + 6 = 10
A) Linear equation B) Equation which is not linear C) Expression

Answer by kietra(57)

3x + 6 = 10. Before you can identify which of the 3 choices it is, you must know what the three choices are. Choice A (linear equation) is any equation that can be written in the form Ax + By = C. An example would be 5x + 5y = 10. An example of an equation that is not a linear equation would be x^2 +y^2 = 9. Choice B (an equation that is nonlinear) would be anything not able to be written in the form Ax+By=C- see examples in choice A. Choice C (an expression) does not have an equal sign and cannot be solved.
Therefore, 3x + 6 = 10 is (A) an example of a linear equation.